Unity Limited acquired 100% of the share capital of Bellvista Limited. Bellvista had issued share capital of $200 000. The book values of Bellvista Limited’s assets were: buildings $100 000, machinery $120 000. The fair values of these assets were: buildings $180 000, machinery $140 000. The tax rate is 30%.

The fair value of the identifiable net assets is:

Select one:

a. $220 000.

b. $270 000.

c. $320 000.

d. $200 000.
Answer #1

Value of Assets:

Building = Fair Value - Book Value = 1,80,000 - 1,00,000 = $80,000

Machinery = Fair Value - Book Value = 1,40,000 - 1,20,000 = $20,000

Acutal Value = Total Value - 70% = (80,000 + 20,000)*70% = $70,000

Fair Value of the identifiable net assets = Share apital + Actual Value = 2,00,000 + 70,000 = $2,70,000.

Hence, The answer is option (b)* = $ 2,70,000